A turbine operating under steady flow conditions receives steam at the





following state : pressure 13.8 bar ; specific volume 0.143 m3/kg ;





internal energy 2590 kJ/kg ; velocity 30 m/s. The state of the steam





leaving the turbine is : pressure 0.35 bar ; specific volume 4.37 m3/kg





; internal energy 2360





kJ/kg ; velocity 90 m/s. Heat is lost to the surroundings at the rate of





0.25 kJ/s. If the rate of steam flow is 0.38 kg/s, what is the power





developed by the turbine ?





1
Expert's answer
2022-02-15T16:54:02-0500

"\dot{m}(u_1+P_1V_1+\frac{C_1^2}{2})-Q=\dot{m}(u_2+P_2V_2+{\frac{C_2^2}{2}})+\dot{w}"



"0.38 \begin{bmatrix}
 2590×10^3+(13.8×10^5)(0.143)+\frac{30}{2}^2 \\
 
\end{bmatrix}-0.25×10^3"

"=1059110.2"


"0.38\begin{bmatrix}
 2360×10^3+(0.35×10^5)(4.37)+\frac{90}{2}^2 \\
 
\end{bmatrix}+\dot{w}"

"=956460+\dot{w}"


"1059110.2=956460+\dot{w}"

"\dot{w}=102650.2"

"=102.65kW"